If we plot of frequency of radiation against the stopping potential of photoelectron, it`s slope would be

Grade:11

1 Answers

Vikas TU
7 years ago
From the Planck's energy relation we get,
hv = hv0 + eVs
dividing by h the whole eqn.
v  = v0 + eVs/h
y = mx + x comparing to this we get,
m = e/h that is the slope of the eqn.